OpenStudy (anonymous):

The sum of the measures of the complement and the supplement of a base angle of an isosceles triangle is 206°. What is the measure of the vertex angle of the triangle?

OpenStudy (anonymous):

26 degree

OpenStudy (anonymous):

hmm thats one of the answer choices that im given

OpenStudy (anonymous):

Let base angles be x so compliment of that angle is 90-x and supplement of that angle is 180-x. So 180-x+90-x=206 360-2x=206 -2x=206-360 -2x=-154 x=-154/-2=77 2x+ vertex angle = 180 154+vertex angle =180 vertex angle = 26

OpenStudy (anonymous):

Oh sorry 180+90 is 270. I typed 360. Sorry

OpenStudy (anonymous):

Thank you heres another question i need help with what is the value of x in the figure below?

OpenStudy (anonymous):